OpenStudy (anonymous):

Solve the system of equations by graphing. {3x+y=-7 {x+2y=-4

OpenStudy (anonymous):

The red line is the first equation

OpenStudy (anonymous):

I dont have a red line

OpenStudy (anonymous):

ok, your colors are different than mine. The first equation is the one that goes higher on the left.

OpenStudy (anonymous):

Ok. I am supposed to graph the equations but I am not sure how to get the ordered pairs to graph them. Can you help me please?

OpenStudy (anonymous):

Sure, that was just the end result to show you what it should look like when you are done. An easy way is to set x= 0 for each equation separately and get the "y". Then set y=0 for each equation separately and get the "x". You'll have 2 sets of two points. Draw a line through them.

OpenStudy (anonymous):

There is a "3-point" solution also and is recommended, because it also algebraically checks your graphical solution at the same time. You solve the equations algebraically to get the one and only one common point, and then set y (or x) equal to zero for each equation separately to get the other variable.

OpenStudy (anonymous):

ok, let me try it. Please stand by....

OpenStudy (anonymous):

For the first equation: 3(0) + y = -7 so one point is (0, -7) 3x + 0 = -7 so the other point is (-7/3, 0)

OpenStudy (anonymous):

Now, you try this method for the second equation and use what I did as a pattern.

OpenStudy (anonymous):

I got it wrong :-( I have one more try. {2x+y=6 {x+4y=-4

OpenStudy (anonymous):

For the second equation: (0) + 2y = -4 so one point is (0, -2) x + 2(0) = -4 so the other point is (-4, 0) Now, for each pair, draw a line through the points.

OpenStudy (anonymous):

Here is your updated graph with tho above 4 points drawn on it.

OpenStudy (anonymous):

And they intersect at (-2, -1)

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

no sorry, still confused!

OpenStudy (anonymous):

so I draw the line at -2,-1?

OpenStudy (anonymous):

I have posts that start out with "For the first equation" and "For the second equation". These are 2 separate posts. In each, I have determined 2 sets of points. You draw a line through the first set of 2 points. Then you go to the other post and get the other points and draw a second line through those points. 2 posts. 2 lines. There is no more work for you to do, because if you just read through my posts, you'll see that I did every last bit of work for you. Nothing left to do.

OpenStudy (anonymous):

ok thanks